OpenStudy (anonymous):

A solution is made by dissolving 25.5 grams of glucose (C6H12O6) in 398 grams of water. What is the freezing point depression of the solvent if the freezing point constant is -1.86 °C/m?

OpenStudy (anonymous):

[delta]T=i Kf m?

OpenStudy (aaronq):

Gotta use the freezing point equation: \(\sf \Delta T=i*m*K_f\) where \(\Delta T\) is the change in temp, \(i\) is the van't hoff constant, m is the molality, and Kf is the depression constant (given in the question). (we can ignore \(i\) because glucose doesnt dissociate into more particles in water). First find the molality though

OpenStudy (aaronq):

yep

OpenStudy (anonymous):

M=number of moles/Volume

OpenStudy (aaronq):

that's molaRity, this is molaLity

OpenStudy (anonymous):

oh my b

OpenStudy (aaronq):

\(\sf molality=\dfrac{moles~of~solute}{kg~of~solution}\)

OpenStudy (aaronq):

lol no worries

OpenStudy (anonymous):

I'm sorry how do i get my 25.5 g into moles?

OpenStudy (anonymous):

25.5g x 1mol / 180g = 0.1417 mol

OpenStudy (aaronq):

\(\sf moles=\dfrac{mass}{Molar~mass}\) yup thats it

OpenStudy (anonymous):

m= 0.1417mol/ 0.398kg= 0.3359m

OpenStudy (anonymous):

0.3559m x -1.86 C/m= 0.662C ??

OpenStudy (aaronq):

the constant has a negative sign, so the answer should be negative

OpenStudy (anonymous):

is that right?

OpenStudy (aaronq):

you need a negative sign though

OpenStudy (aaronq):

otherwise it's good

OpenStudy (anonymous):

-0.662 C

OpenStudy (aaronq):

yeah thats it

OpenStudy (anonymous):

Thanks!

OpenStudy (aaronq):

np!